Reymond has a job with starting salary of 14,000. Suppose he receives a yearly increase of 2,500, how much would his salary be on his fifth year on the job?
Answer by macston(5194)

.
For fifth year on the job:
Salary=14000+2500(year on the job-1)
Salary=14000+(2500(5-1))
Salary=14000+2500(4)
Salary=14000+10000
Salary=24000